# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 4.1 Examples
We show here that, by following an approach similar to that used in Example 27, one can also find a Darboux transformation for the Camassa-Holm =$\mathcal{E}=\{z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t}\}$. To this end one can consider the $2$-dimensional differentiable covering of $\mathcal{E}$ defined now by
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\hat{\rho}{}_{x}=\frac{1}{2\hat{\eta}}\hat{\rho}^{2}+\frac{z_{xx}}{2}+\frac{-{\hat{\eta}}^{2}-\hat{\eta}\,z}{2\,\hat{\eta}},\vspace{10pt}\\ \hat{\rho}{}_{t}=-\frac{\left(-2\,\hat{\eta}+z\right)}{2\,\hat{\eta}}\hat{\rho}^{2}-z{}_{x}\,\hat{\rho}-\frac{z\,z{}_{xx}}{2}-\frac{2\,\hat{\eta}^{3}+\hat{\eta}^{2}z-\hat{\eta}\,z{}^{2}}{2\,\hat{\eta}},\vspace{10pt}\\ \rho_{x}=\frac{1}{2\eta}\rho^{2}+\frac{z_{xx}}{2}+\frac{-{\eta}^{2}-\eta\,z}{2\,\eta},\vspace{10pt}\\ \rho_{t}=-\frac{\left(-2\,\eta+z\right)}{2\,\eta}\rho^{2}-z_{x}\,\rho-\frac{z\,z_{xx}}{2}-\frac{2\,\eta^{3}+\eta^{2}z-\eta\,z^{2}}{2\,\eta},\vspace{10pt}\\ z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t},\end{array}\right.$
which is obtained by doubling the Riccati system (35). Hence, being (36) still a conservation law for $\mathcal{Y}$, one can search for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
@@
relatively to $c\,\mu$, with $c\in\mathbb{R}$. In this case, when $\hat{\eta}\neq\eta$ and $\hat{\rho}\neq\rho$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=-\partial_{x}+(2\rho-z_{x})\,\partial_{z}+\frac{\left(\eta^{2}-\rho^{2}\right)}{2\eta}\partial_{\rho}+\frac{\eta^{2}\hat{\eta}-\eta\,\hat{\eta}^{2}-2\eta\,\rho\,\hat{\rho}+\hat{\eta}\,\rho^{2}+\eta\,\hat{\rho}^{2}}{2\,\eta\,\left(\eta-\hat{\eta}\right)}\partial_{\hat{\rho}}+...,$
$\begin{array}[]{l}\xi^{1}=x-\ln\left(\frac{\rho-\eta}{\rho+\eta}\right),\quad\xi^{2}=t,\quad\nu^{1}=\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)z+\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}\,z_{x},\vspace{10pt}\\ \nu^{2}=\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}z+\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)\,z_{x},\quad\nu^{3}=\frac{\hat{\eta}\,\rho^{2}-\hat{\eta}\,\rho\hat{\rho}-\eta\,\hat{\eta}\,\left(\eta-\hat{\eta}\right)}{\eta\,\hat{\rho}-\hat{\eta}\,\rho}\end{array}$
a basic system of first-order order invariants of $Y$, one finds that in this case
@@
$\mathcal{Y}^{\prime}:\;\left\{\begin{array}[]{l}\rho^{\prime}_{x^{\prime}}=\frac{1}{2\eta}\rho^{\prime}{}^{2}+\frac{z^{\prime}_{x^{\prime}x^{\prime}}}{2}+\frac{-{\eta}^{2}-\eta\,z^{\prime}}{2\,\eta},\vspace{10pt}\\ \rho^{\prime}_{t^{\prime}}=-\frac{\left(-2\,\eta+z^{\prime}\right)}{2\,\eta}\rho^{\prime}{}^{2}-z^{\prime}_{x^{\prime}}\,\rho^{\prime}-\frac{z^{\prime}\,z^{\prime}_{x^{\prime}x^{\prime}}}{2}-\frac{2\,\eta^{3}+\eta^{2}z^{\prime}-\eta\,z^{\prime}{}^{2}}{2\,\eta},\vspace{10pt}\\ z^{\prime}_{t^{\prime}x^{\prime}x^{\prime}}=3\,z^{\prime}\,z^{\prime}_{x^{\prime}}-z^{\prime}\,z^{\prime}_{x^{\prime}x^{\prime}x^{\prime}}-2\,z^{\prime}_{x^{\prime}}\,z^{\prime}_{x^{\prime}x^{\prime}}+z^{\prime}_{t^{\prime}}.\end{array}\right.$
@@
We consider here the Harry-Dym equation =$\mathcal{E}=\{z_{t}=z^{3}z_{xxx}\}$ and show how the invariants of a nonlocal pseudosymmetry $Y$ can be used to determine an ABT of this equation. To this end we consider the following ZCR
$\alpha=\left(\begin{array}[]{ll}0&4\eta\\ -\frac{\eta}{z^{2}}&0\end{array}\right)dx+\left(\begin{array}[]{ll}8\eta^{2}z_{x}&-64\eta^{3}z\\ 2\eta z_{xx}+\frac{16\eta^{3}}{z}&-8\eta^{2}z_{x}\end{array}\right)dt,\qquad\eta\in\mathbb{R},.$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}=-4\eta\rho^{2}-\frac{1}{z^{2}}\vspace{5pt}\\ \rho_{t}=64\eta^{3}\,z\,\rho^{2}-16\eta^{2}z_{x}\rho+2\eta z_{xx}+\frac{16\,\eta^{3}}{z}\vspace{5pt}\\ z_{t}=z^{3}z_{xxx}\end{array}\right.$
@@
First, by searching for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=\frac{1}{8\eta}\partial_{x}+z\,\rho\,\partial_{z}-\frac{\rho^{2}}{2}\,\partial_{\rho}+...,$
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x-1/(4\eta\rho),\,t,\,z\,R^{2}\}$ a basic system of zeroth order invariants of $Y$, one is naturally lead to consider $\xi^{1}=x-1/(4\eta\rho),\,\xi^{2}=t$ and $\nu=\nu\left(z\,R^{2}\right)$. This way, a direct computation shows that $z^{\prime}_{t^{\prime}}-z^{\prime}{}^{3}z^{\prime}_{x^{\prime}x^{\prime}x^{\prime}}=0\mod{\mathcal{E}^{(\infty)}}$ iff $\nu=-1/(4z\rho^{2})$. Hence, one has the ABT of $\mathcal{E}$
@@
provided by the invariants of above nonlocal pseudosymmetry $Y$ of $\mathcal{E}$.
@@
We consider here the coupled Schrodinger system =$\mathcal{E}=\{z_{t}^{1}=2\left(z^{1}\right)^{2}z^{2}+z_{xx}^{1},\,z_{t}^{2}=-2z^{1}\left(z^{2}\right)^{2}-z_{xx}^{2}\}$ and show how the invariants of a nonlocal pseudosymmetry $Y$ can be used to determine an ABT of $\mathcal{E}$. To this end we consider the following ZCR
$\alpha=\left(\begin{array}[]{cc}\eta&-z^{2}\vspace{5pt}\\ z^{1}&-\eta\end{array}\right)dx+\left(\begin{array}[]{cc}-2\eta^{2}-z^{1}\,z^{2}&2\eta z^{2}+z_{x}^{2}\vspace{5pt}\\ -2\eta z^{1}+z_{x}^{1}&2\eta^{2}+z^{1}z^{2}\end{array}\right)dt,\qquad\eta\in\mathbb{R},.$
@@
$\left\{\begin{array}[]{l}\rho_{x}=z^{2}\rho^{2}-2\eta\rho+z^{1}\vspace{5pt}\\ \rho_{t}=\left(-2\eta\,z^{2}-z_{x}^{2}\right)\rho^{2}+\left(2z^{1}z^{2}+4\eta^{2}\right)\rho-2\eta z^{1}+z_{x}^{1}\vspace{5pt}\\ z_{t}^{1}=2\left(z^{1}\right)^{2}z^{2}+z_{xx}^{1},\vspace{5pt}\\ \,z_{t}^{2}=-2z^{1}\left(z^{2}\right)^{2}-z_{xx}^{2},\end{array}\right.$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\hat{\rho}_{x}=z^{2}\hat{\rho}^{2}-2\hat{\eta}\hat{\rho}+z^{1}\vspace{5pt}\\ \hat{\rho}_{t}=\left(-2\hat{\eta}\,z^{2}-z_{x}^{2}\right)\hat{\rho}^{2}+\left(2z^{1}z^{2}+4\hat{\eta}^{2}\right)\hat{\rho}-2\hat{\eta}z^{1}+z_{x}^{1}\vspace{5pt}\\ \rho_{x}=z^{2}\rho^{2}-2\eta\rho+z^{1}\vspace{5pt}\\ \rho_{t}=\left(-2\eta\,z^{2}-z_{x}^{2}\right)\rho^{2}+\left(2z^{1}z^{2}+4\eta^{2}\right)\rho-2\eta z^{1}+z_{x}^{1}\vspace{5pt}\\ z_{t}^{1}=2\left(z^{1}\right)^{2}z^{2}+z_{xx}^{1},\vspace{5pt}\\ \,z_{t}^{2}=-2z^{1}\left(z^{2}\right)^{2}-z_{xx}^{2}.\end{array}\right.$
Hence, being $\mu$ still a conservation law for $\mathcal{Y}$, one can search for pseudosymmetries of $\mathcal{Y}$ that are pseudo-prolongations of a vector field
$\begin{array}[]{l}a_{1}(x,t,z^{1},z^{2},\rho,\hat{\rho})\partial_{x}+a_{2}(x,t,z^{1},z^{2},\rho,\hat{\rho})\partial_{t}+b^{11}(x,t,z^{1},z^{2},\rho,\hat{\rho})\partial_{z^{1}}+b^{12}(x,t,z^{1},z^{2},\rho,\hat{\rho})\partial_{z^{2}}\vspace{10pt}\\ \qquad+b^{21}(x,t,z^{1},z^{2},\rho,\hat{\rho})\partial_{\rho},+b^{22}(x,t,z^{1},z^{2},\rho,\hat{\rho})\partial_{\hat{\rho}}\;,\end{array}$
relatively to $c\,\mu$, with $c\in\mathbb{R}$. In this case, when $\hat{\eta}\neq\eta$ and $\hat{\rho}\neq\rho$, one finds that the most general pseudosymmetry of this type is a constant multiple of
$Y=\rho^{2}\,\partial_{z^{1}}-\partial_{z^{2}}-\frac{\left(\rho-\hat{\rho}\right)^{2}}{2(\eta-\hat{\eta})}\partial_{\hat{\rho}}+...,$
@@
$\xi^{1}=x,\quad\xi^{2}=t,\quad\nu^{1}=\frac{((2\eta-2\hat{\eta})\rho\hat{\rho}+z^{1}\left(\rho-\hat{\rho}\right)}{\rho-\hat{\rho}},\quad\nu^{2}=\frac{(\rho-\hat{\rho})z^{2}-2\eta+2\hat{\eta}}{\rho-\hat{\rho}},\quad\nu^{3}=\rho$
a basic system of zeroth-order invariants of $Y$, one finds that in this case
@@
We consider here Tzitzeica equation =$\mathcal{E}=\{z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}\}$. As already shown in Example 25, $\mathcal{E}$ admits the nonlocal conservation law $\mu=\lambda{\small\rho^{2}}\,dx+\frac{1}{\lambda}e^{z}{\small\rho^{1}}\,dt$ in the $1$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}^{1}=-\lambda\,\rho^{1}\rho^{2}-z_{x}\,\rho^{1}+\lambda,\vspace{5pt}\\ \rho_{t}^{1}=-{{\rm e}^{z}}{\left(\rho^{1}\right)}^{2}+\frac{1}{\lambda}e^{-2z}\rho^{1},\vspace{5pt}\\ \rho_{x}^{2}=-\lambda{\left(\rho^{2}\right)}^{2}+z_{x}{\rho^{2}}+\lambda\rho^{1},\vspace{5pt}\\ \rho_{t}^{2}=-\frac{1}{\lambda}{{\rm e}^{z}}{\rho^{1}}{\rho^{2}}+\frac{1}{\lambda}e^{z},\vspace{5pt}\\ z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}.\end{array}\right.$
@@
Contrary to previous examples, in this case, we did not find any nontrivial pseudosymmetry, which would be a pseudo-prolongation of some vector field relative to a multiple of $\mu$. However, we found that $\mathcal{Y}$ admits a $2$-pseudosymmetry
@@
$\mathbb{Y}=\left[\partial_{z}-\frac{2\rho^{1}\rho^{2}-1}{2\rho^{1}}\partial_{\rho^{2}}+\ldots\;,\quad\partial_{\rho^{1}}-\frac{\rho^{2}}{\rho^{1}}\partial_{\rho^{2}}+\ldots\right]^{T},$
given by pseudo-prolongations relative to the following horizontal $1$-form
$\begin{array}[]{l}\gamma=\left(\begin{array}[]{cc}-{\displaystyle\frac{4\lambda\,(\rho^{1}\rho^{2})^{2}+2\lambda\,(\rho^{1})^{3}-4\lambda\,\rho^{1}\rho^{2}+\lambda\,}{2\rho^{2}(\rho^{1})^{2}-\rho^{1}}\vspace{10pt}}&{\displaystyle\frac{12\lambda\,(\rho^{1}\rho^{2})^{2}+4\lambda\,(\rho^{1})^{3}-12\lambda\,\rho^{1}\rho^{2}+3\lambda}{4\rho^{1}\rho^{2}-2}}\\ -{\displaystyle\frac{4\lambda\,\rho^{1}(\rho^{2})^{2}+4\lambda\,(\rho^{1})^{2}-2\lambda\,\rho^{2}}{2\rho^{2}(\rho^{1})^{2}-\rho^{1}}}&{\displaystyle\frac{4\lambda(\rho^{1})^{2}+4\lambda\rho^{1}(\rho^{2})^{2}-2z_{x}\rho^{1}\rho^{2}-2\lambda\rho^{2}+z_{x}}{2\rho^{1}\rho^{2}-1}}\end{array}\right)dx\vspace{15pt}\\ \qquad+\left(\begin{array}[]{cc}{\displaystyle\frac{4{\rm e}^{z}(\rho^{1})^{3}\rho^{2}-2{\rm e}^{z}(\rho^{1})^{2}+6\rho^{1}\,(\rho^{2})^{2}{\rm e}^{-2\,z}-2\rho^{2}{\rm e}^{-2\,z}}{\lambda\,\left(2\rho^{2}(\rho^{1})^{2}-\rho^{1}\right)}\vspace{10pt}}&-{\displaystyle\frac{{\rm e}^{z}(\rho^{1})^{2}}{\lambda}-3\frac{\rho^{2}{\rm e}^{-2z}}{\lambda}+\frac{{\rm e}^{-2z}}{2\lambda\rho^{1}}}\\ {\displaystyle\frac{4{\rm e}^{z}(\rho^{1})^{2}\rho^{2}-2\,{\rm e}^{z}\rho^{1}+4(\rho^{2})^{2}{\rm e}^{-2z}}{\lambda\,\left(2\rho^{2}(\rho^{1})^{2}-\rho^{1}\right)}}&-{\displaystyle 2\frac{{\rm e}^{3z}\rho^{1}}{\lambda}-\frac{\rho^{2}\,{\rm e}^{-2z}}{\lambda\rho^{1}}}\end{array}\right)dt,\end{array}$
@@
that satisfies the zero-curvature condition $d_{H}\mu-\mu\wedge\mu=0\mod{\mathcal{E}^{(\infty)}}$. Such a $2$-pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal $2$-pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$.

### v5

@@ 4.1 Examples
We show here that, by following an approach similar to that used in Example 28, one can also find a Darboux transformation for the Camassa-Holm $\mathcal{E}=\{z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t}\}$. To this end one can consider the $2$-dimensional differentiable covering of $\mathcal{E}$ defined now by
$\mathcal{Y}:\;\left\{{\displaystyle\begin{array}[]{l}{\displaystyle\hat{\rho}{}_{x}=\frac{1}{2\hat{\eta}}\hat{\rho}^{2}+\frac{z_{xx}}{2}+\frac{-{\hat{\eta}}^{2}-\hat{\eta}\,z}{2\,\hat{\eta}}},\vskip 10.0pt\\ {\displaystyle\hat{\rho}{}_{t}=-\frac{\left(-2\,\hat{\eta}+z\right)}{2\,\hat{\eta}}\hat{\rho}^{2}-z{}_{x}\,\hat{\rho}-\frac{z\,z{}_{xx}}{2}-\frac{2\,\hat{\eta}^{3}+\hat{\eta}^{2}z-\hat{\eta}\,z{}^{2}}{2\,\hat{\eta}}},\vskip 10.0pt\\ {\displaystyle\rho_{x}=\frac{1}{2\eta}\rho^{2}+\frac{z_{xx}}{2}+\frac{-{\eta}^{2}-\eta\,z}{2\,\eta}},\vskip 10.0pt\\ {\displaystyle\rho_{t}=-\frac{\left(-2\,\eta+z\right)}{2\,\eta}\rho^{2}-z_{x}\,\rho-\frac{z\,z_{xx}}{2}-\frac{2\,\eta^{3}+\eta^{2}z-\eta\,z^{2}}{2\,\eta}},\vskip 10.0pt\\ z_{txx}=3\,z\,z_{x}-z\,z_{xxx}-2\,z_{x}\,z_{xx}+z_{t},\end{array}}\right$.
which is obtained by doubling the Riccati-type system (41). Hence, being (42) still a conservation law for the new system $\mathcal{Y}$, one can search for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of (relative) vector fields of the form
@@
relatively to $2\,\mu$. In this case, when $\hat{\eta}\neq\eta$ and $\hat{\rho}\neq\rho$, one finds that the most general pseudosymmetry of this type with $b^{1}\neq 0$ is a constant multiple of
$\bar{Y}={\displaystyle-\partial_{x}+(2\rho-z_{x})\,\partial_{z}+\frac{\left(\eta^{2}-\rho^{2}\right)}{2\eta}\partial_{\rho}+\frac{\eta^{2}\hat{\eta}-\eta\,\hat{\eta}^{2}-2\eta\,\rho\,\hat{\rho}+\hat{\eta}\,\rho^{2}+\eta\,\hat{\rho}^{2}}{2\,\eta\,\left(\eta-\hat{\eta}\right)}\partial_{\hat{\rho}}+\ldots\,.}$
Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being
$\begin{array}[]{l}{\displaystyle\xi^{1}=x-\ln\left(\frac{\rho-\eta}{\rho+\eta}\right),\quad\xi^{2}=t,\quad\nu^{1}=\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)z+\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}\,z_{x},}\vskip 10.0pt\\ {\displaystyle\nu^{2}=\frac{2\,\eta\,\rho}{\rho^{2}-\eta^{2}}z+\left(\frac{\rho^{2}+\eta^{2}}{\rho^{2}-\eta^{2}}\right)\,z_{x},\quad\nu^{3}=\frac{\hat{\eta}\,\rho^{2}-\hat{\eta}\,\rho\hat{\rho}-\eta\,\hat{\eta}\,\left(\eta-\hat{\eta}\right)}{\eta\,\hat{\rho}-\hat{\eta}\,\rho},}\end{array}$
a basic system of first-order order invariants of $\bar{Y}$, one finds that in this case
@@
$\mathcal{Y}^{\prime}:\;\left\{\begin{array}[]{l}{\displaystyle\rho^{\prime}_{x^{\prime}}=\frac{1}{2\eta}\rho^{\prime}{}^{2}+\frac{z^{\prime}_{x^{\prime}x^{\prime}}}{2}+\frac{-{\eta}^{2}-\eta\,z^{\prime}}{2\,\eta},}\vskip 10.0pt\\ {\displaystyle\rho^{\prime}_{t^{\prime}}=-\frac{\left(-2\,\eta+z^{\prime}\right)}{2\,\eta}\rho^{\prime}{}^{2}-z^{\prime}_{x^{\prime}}\,\rho^{\prime}-\frac{z^{\prime}\,z^{\prime}_{x^{\prime}x^{\prime}}}{2}-\frac{2\,\eta^{3}+\eta^{2}z^{\prime}-\eta\,z^{\prime}{}^{2}}{2\,\eta}},\vskip 10.0pt\\ z^{\prime}_{t^{\prime}x^{\prime}x^{\prime}}=3\,z^{\prime}\,z^{\prime}_{x^{\prime}}-z^{\prime}\,z^{\prime}_{x^{\prime}x^{\prime}x^{\prime}}-2\,z^{\prime}_{x^{\prime}}\,z^{\prime}_{x^{\prime}x^{\prime}}+z^{\prime}_{t^{\prime}}.\end{array}\right$.
@@
We consider here the Harry-Dym equation $\mathcal{E}=\{z_{t}=z^{3}z_{xxx}\}$ and show how the invariants of a nonlocal pseudosymmetry $\bar{Y}$ can be used to determine an auto-Bäcklund transformation of this equation. To this end we consider the following ZCR
$\alpha=\left({\displaystyle\begin{array}[]{ll}0&4\eta\\ -\frac{\eta}{z^{2}}&0\end{array}}\right)dx+\left({\displaystyle\begin{array}[]{ll}8\eta^{2}z_{x}&-64\eta^{3}z\\ 2\eta z_{xx}+\frac{16\eta^{3}}{z}&-8\eta^{2}z_{x}\end{array}}\right)dt,\qquad\eta\in\mathbb{R}-\{0\}.$
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}{\displaystyle\rho_{x}=-4\eta\rho^{2}-\frac{1}{z^{2}}},\vskip 5.0pt\\ {\displaystyle\rho_{t}=64\eta^{3}\,z\,\rho^{2}-16\eta^{2}z_{x}\rho+2\eta z_{xx}+\frac{16\,\eta^{3}}{z}},\vskip 5.0pt\\ z_{t}=z^{3}z_{xxx},\end{array}\right$.
@@
First, by searching for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of vector fields of the form
@@
relatively to conservation laws $c\,\mu$, with $c\in\mathbb{R}-\{0\}$, one finds that the most general pseudosymmetry of this type with $b^{1}\neq 0$ is a constant multiple of
$\bar{Y}=\frac{1}{8\eta}\partial_{x}+z\,\rho\,\partial_{z}-\frac{\rho^{2}}{2}\,\partial_{\rho}+\ldots\,$,
with $c=2$. Such a pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$. Thus, being $\{x-1/(4\eta\rho),\,t,\,z\,\rho^{2}\}$ a basic system of zeroth order invariants of $\bar{Y}$, one is naturally lead to consider $\xi^{1}=x-1/(4\eta\rho),\,\xi^{2}=t$ and taking $\nu=-1/(4z\rho^{2})$ one obtains $z^{\prime}_{t^{\prime}}-z^{\prime}{}^{3}z^{\prime}_{x^{\prime}x^{\prime}x^{\prime}}=0$. Hence, one has the auto-Bäcklund transformation of $\mathcal{E}$
@@
provided by the invariants of the pseudosymmetry $\bar{Y}$.
@@
We consider here the coupled Schrödinger system $\mathcal{E}=\{z_{t}^{1}=2\left(z^{1}\right)^{2}z^{2}+z_{xx}^{1},\,z_{t}^{2}=-2z^{1}\left(z^{2}\right)^{2}-z_{xx}^{2}\}$ and show how the invariants of a nonlocal pseudosymmetry $\bar{Y}$ can be used to determine an auto-Bäcklund transformation of $\mathcal{E}$. To this end we consider the following ZCR
$\alpha=\left(\begin{array}[]{cc}\eta&-z^{2}\vskip 5.0pt\\ z^{1}&-\eta\end{array}\right)dx+\left(\begin{array}[]{cc}-2\eta^{2}-z^{1}\,z^{2}&2\eta z^{2}+z_{x}^{2}\vskip 5.0pt\\ -2\eta z^{1}+z_{x}^{1}&2\eta^{2}+z^{1}z^{2}\end{array}\right)dt,\qquad\eta\in\mathbb{R},.$
@@
$\left\{\begin{array}[]{l}\rho_{x}=z^{2}\rho^{2}-2\eta\rho+z^{1},\vskip 5.0pt\\ \rho_{t}=\left(-2\eta\,z^{2}-z_{x}^{2}\right)\rho^{2}+\left(2z^{1}z^{2}+4\eta^{2}\right)\rho-2\eta z^{1}+z_{x}^{1},\vskip 5.0pt\\ z_{t}^{1}=2\left(z^{1}\right)^{2}z^{2}+z_{xx}^{1},\vskip 5.0pt\\ \,z_{t}^{2}=-2z^{1}\left(z^{2}\right)^{2}-z_{xx}^{2},\end{array}\right$.
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\hat{\rho}_{x}=z^{2}\hat{\rho}^{2}-2\hat{\eta}\hat{\rho}+z^{1},\vskip 5.0pt\\ \hat{\rho}_{t}=\left(-2\hat{\eta}\,z^{2}-z_{x}^{2}\right)\hat{\rho}^{2}+\left(2z^{1}z^{2}+4\hat{\eta}^{2}\right)\hat{\rho}-2\hat{\eta}z^{1}+z_{x}^{1},\vskip 5.0pt\\ \rho_{x}=z^{2}\rho^{2}-2\eta\rho+z^{1},\vskip 5.0pt\\ \rho_{t}=\left(-2\eta\,z^{2}-z_{x}^{2}\right)\rho^{2}+\left(2z^{1}z^{2}+4\eta^{2}\right)\rho-2\eta z^{1}+z_{x}^{1},\vskip 5.0pt\\ z_{t}^{1}=2\left(z^{1}\right)^{2}z^{2}+z_{xx}^{1},\vskip 5.0pt\\ \,z_{t}^{2}=-2z^{1}\left(z^{2}\right)^{2}-z_{xx}^{2}.\end{array}\right$.
Hence, being $\mu$ still a conservation law for $\mathcal{Y}$, one can search for pseudosymmetries $\bar{Y}:=\left.Y\right|_{\mathcal{Y}^{(\infty)}}$ of $\mathcal{Y}$ determined by the pseudoprolongations $Y$ of vector fields of the form
[…]
relatively to $c\,\mu$, with $c\in\mathbb{R}-\{0\}$. In this case, when $\hat{\eta}\neq\eta$ and $\hat{\rho}\neq\rho$, one finds that the most general pseudosymmetry of this type with $b^{11}b^{12}\neq 0$ is a constant multiple of
$\bar{Y}=\rho^{2}\,\partial_{z^{1}}-\partial_{z^{2}}-\frac{\left(\rho-\hat{\rho}\right)^{2}}{2(\eta-\hat{\eta})}\partial_{\hat{\rho}}$$\ldots\,$,
@@
$\xi^{1}=x,\quad\xi^{2}=t,\quad\nu^{1}=\frac{((2\eta-2\hat{\eta})\rho\hat{\rho}+z^{1}\left(\rho-\hat{\rho}\right)}{\rho-\hat{\rho}},\quad\nu^{2}=\frac{(\rho-\hat{\rho})z^{2}-2\eta+2\hat{\eta}}{\rho-\hat{\rho}},\quad\nu^{3}=\rho,$
a basic system of zeroth-order invariants of $\bar{Y}$, one finds that in this case
@@
We consider here Tzitzeica equation $\mathcal{E}=\{z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}\}$. As already shown in Example 26, $\mathcal{E}$ admits the nonlocal conservation law $\mu=\lambda{\small\rho^{2}}\,dx+\frac{1}{\lambda}e^{z}{\small\rho^{1}}\,dt$ in the $1$-dimensional differentiable covering $\mathcal{Y}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$ defined by
@@
$\mathcal{Y}:\;\left\{\begin{array}[]{l}\rho_{x}^{1}=-\lambda\,\rho^{1}\rho^{2}-z_{x}\,\rho^{1}+\lambda,\vskip 5.0pt\\ {\displaystyle\rho_{t}^{1}=-\frac{1}{\lambda}{{\rm e}^{z}}{\left(\rho^{1}\right)}^{2}+\frac{1}{\lambda}e^{-2z}\rho^{2}},\vskip 5.0pt\\ \rho_{x}^{2}=-\lambda{\left(\rho^{2}\right)}^{2}+z_{x}{\rho^{2}}+\lambda\rho^{1},\vskip 5.0pt\\ {\displaystyle\rho_{t}^{2}=-\frac{1}{\lambda}{{\rm e}^{z}}{\rho^{1}}{\rho^{2}}+\frac{1}{\lambda}e^{z}},\vskip 5.0pt\\ {\displaystyle z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}}.\end{array}\right$.
@@
Contrary to previous examples, in this case, we did not find any nontrivial pseudosymmetry, which would be a pseudoprolongation of some vector field relative to a multiple of $\mu$. However, we found that $\mathcal{Y}$ admits a $2$-pseudosymmetry given by
@@
$\bar{\mathbb{Y}}=\left.\mathbb{Y}\right|_{\mathcal{Y}^{(\infty)}}=\left[\partial_{z}-\frac{2\rho^{1}\rho^{2}-1}{2\rho^{1}}\partial_{\rho^{2}}+\ldots\;,\quad\partial_{\rho^{1}}-\frac{\rho^{2}}{\rho^{1}}\partial_{\rho^{2}}+\ldots\right]^{T},$
induced on $\mathcal{Y}^{(\infty)}$ by a pseudoprolongation $\mathbb{Y}$ relative to the horizontal $1$-form
γ=(4λ(ρ1ρ2)2+2λ(ρ1)34λρ1ρ2+λ2ρ2(ρ1)2ρ112λ(ρ1ρ2)2+4λ(ρ1)312λρ1ρ2+3λ4ρ1ρ224λρ1(ρ2)2+4λ(ρ1)22λρ22ρ2(ρ1)2ρ14λ(ρ1)2+4λρ1(ρ2)22zxρ1ρ22λρ2+zx2ρ1ρ21)dx
+(4ez(ρ1)3ρ22ez(ρ1)2+6ρ1(ρ2)2e2z2ρ2e2zλ(2ρ2(ρ1)2ρ1)ez(ρ1)2λ3ρ2e2zλ+e2z2λρ14ez(ρ1)2ρ22ezρ1+4(ρ2)2e2zλ(2ρ2(ρ1)2ρ1)2ezρ1λρ2e2zλρ1)dt,
[…]
@@
satisfying the zero-curvature condition $d_{H}\gamma-\gamma\wedge\gamma=0\mod{\mathcal{E}^{(\infty)}}$. Such a $2$-pseudosymmetry of $\mathcal{Y}$ can be seen as a nonlocal $2$-pseudosymmetry of $\mathcal{E}$, with respect to the differentiable covering defined by $\mathcal{Y}$.